Three block of mass 1kg,4kg,2kg ate placed on smoOth horizontal plane find.  1 - acceleration of the system  2- the normal force between 1 ka  block and 4 ka block 3-  the net force on 2 kg block

Grade:12

2 Answers

Arun
4 years ago
Dear student 
 
Question is not properly written or does not have proper meaning. Please repost the question. You can also attach an attachment.
I will be happy to help you. 
Vikas TU
4 years ago
Dear student 
i).Acceleration = Fnet / mass 
=> (120-50)/1+4+2 = 70/ = 10m/sec^2
ii). F = ma 
120-N = 10 
So, N = 110 N 
iii).F = ma 
=> 2*10 = 20N